# Object Re-identification via Spatial-temporal Fusion Networks and Causal Identity Matching

### v1

@@ Abstract
Object re-identification (ReID) in large camera networks has many challenges. First, the similar appearances of objects degrade ReID performances. This challenge cannot be addressed by existing appearance-based ReID methods. Second, most ReID studies are performed in laboratory settings and do not consider ReID problems in real-world scenarios. To overcome these challenges, we introduce a novel ReID framework that leverages a spatial-temporal fusion network and causal identity matching (CIM). The framework estimates camera network topology using the proposed adaptive Parzen window and combines appearance features with spatial-temporal cue within the Fusion Network. It achieved outstanding performance across several datasets, including VeRi776, Vehicle-3I, and Market-1501, achieving up to 99.70% rank-1 accuracy and 95.5% mAP. Furthermore, the proposed CIM approach, which dynamically assigns gallery sets based on the camera network topology, further improved ReID accuracy and robustness in real-world settings, evidenced by a 94.95% mAP and 95.19% F1 score on the Vehicle-3I dataset. The experimental results support the effectiveness of incorporating spatial-temporal information and CIM for real-world ReID scenarios regardless of the data domain (e.g., vehicle, person).
@@ I Introduction
An increasing number of surveillance cameras have been installed in public places for safety, traffic monitoring, and security. However, monitoring all cameras entails significant human effort and resources. Re-identification (ReID) can be applied to automatically track targets across multiple non-overlapping cameras to reduce human effort. In general, most studies have focused on the visual appearance of a target to perform ReID. For example, many studies [1, 2] have proposed feature learning methods to represent the appearance of a target. Similarly, to measure feature distance between queries and gallery images effectively, metric learning methods [3, 4] have also been proposed. Recently, the ReID performance has been improved by learning visual features and distance metrics with developments in deep learning [5, 6]. These appearance-based methods are robust to pose variations, viewpoint changes, and illumination changes.
(a) Appearance ambiguity in ReID
(b) High computational complexity
Figure 1: Challenges in object re-identification
@@
Nevertheless, appearance ambiguity caused by objects with similar appearances is still not alleviated. As shown in Fig. 1a, vehicles are difficult to match exactly because of their similar appearance due to the same model types. Compared to vehicles, people have relatively distinct features but have similar appearances due to the same clothes. In addition, the appearance of a large number of objects in multiple cameras causes higher computational complexity and lower identification performance because the number of objects with similar appearance to the target increases (Fig. 1b). Thus, relying only on the target appearance is ineffective for the object ReID problem.
@@
Recently, ReID studies using additional spatial and temporal information to alleviate appearance ambiguity [7, 8, 9, 10] have been proposed. Researchers have built a camera-network topology that explains the spatial and temporal relationships between cameras and used the topology to reduce redundant retrieval time ranges for queries. While the methods can potentially improve appearance-based ReID models [5, 6], they still have limitations. Their camera network topology modeling approaches are too simple, and integrating the appearance model with spatial-temporal information lacks optimization.
In addition, most ReID studies have not considered applying ReID in real-world scenarios. For example, videos continuously come in real-time from each camera. However, most ReID studies [5, 11] have performed ReID on fixed and refined query–gallery sets, which are very clean and easy to handle. These laboratory settings are difficult to implement in real-world environments and require additional processes such as object appearance management and query-gallery sets management. In summary, previous studies did not fully consider spatial-temporal dependencies between cameras and real-world scenarios.
@@
To consider the spatial and temporal dependencies between cameras, we proposed a spatial-temporal fusion network (FusionNet) in our previous study [12]. For training FusionNet, we proposed a novel adaptive Parzen window method robust to outliers and sparse responses between camera pairs (Section III-A). It can effectively manage different connection strengths of camera pairs for reliable camera network topology estimation. After estimating the topology, we trained a FusionNet to optimally combine appearance similarity and spatial-temporal probabilities (Section III-B). In this study, we extend our study[12] to explore ReID problems in real-world scenarios extensively. To this end, we first build a new large-scale vehicle ReID dataset called Vehicle-3I, including 2,038 identities captured by 11 synchronized cameras installed in three different intersections (Section V). Unlike many public datasets [13, 14] that provide only cropped image patches, it provides full-frame videos.
@@
Furthermore, we focus on critical and practical issues for ReID in real-world scenarios. Issue 1) Each identity has a set of multiple appearances in the video. Therefore, an appearance management process and ID-to-ID re-identification should be addressed. Issue 2) The query and gallery cannot be predefined as in previous studies. Instead, the query and gallery are determined dynamically based on the target objects (query), and their ReID results within the camera network. In Section IV, we propose the appearance management and ID-to-ID ReID methods by considering multiple appearances of each identity in the video. Also, we perform Causal Identity Matching (CIM), which dynamically controls the query and gallery sets according to the camera connection causality.
@@
To evaluate the proposed spatial-temporal fusion network, we tested the VeRi776 [14], Vehicle-3I vehicle ReID datasets and Market-1501 [13] person ReID dataset. We evaluated the effectiveness of the adaptive Parzen window and FusionNet, and the proposed spatial-temporal fusion network achieved the best performances in all datasets for the rank-1 accuracy (99.70%, 89.22%, 99.11%) and mean average precision (mAP) (91.71%, 83.62%, 93.80%). The results support that our methods can significantly improve the re-id performance regardless of the data domain (e.g., vehicle, person). We additionally perform CIM and evaluate the proposed methods for real-world scenarios; the proposed methods achieved the superior performances in Vehicle-3I dataset for the 94.95% mAP and 95.19% $F_{1}$ score.
@@ II-A Appearance-based ReID
Most ReID studies have focused on developing visual representations of images to distinguish their appearance. There has been extensive research on metric learning and feature learning methods to achieve this objective. For metric learning, the widely studied approach has been learning the Mahalanobis distance [3, 4]. In particular, the optimization of triplet loss for deep metric learning [15, 16, 5] has shown outstanding performance in ReID tasks. Ghosh et al. [6] have introduced Relation Preserving Triplet Mining (RPTM), a scheme for feature matching-guided triplet mining that ensures triplets preserve natural subgroupings in object IDs.
Ahmed et al. [1] initially employed a deep convolutional neural network (CNN) architecture to capture local relationships between two input images based on mid-level features. Chen et al. [2] improved the CNN-based ReID by proposing a deep pyramid feature learning CNN, which is capable of learning scale-specific discriminative features. Similarly, efforts have been made by studies [15, 17] to extract robust local features. To capture more appearance details, Khamis et al. [18] utilized the attributes of targets for ReID. Recently, He et al. [11] proposed TransReID, which is the first attempt to use a transformer to learn robust features from the image patches and Chen et al. [19] utilized Swin Transformer [20] as a backbone for downstream ReID tasks.
CLIP-ReID, which fine-tunes the initialized visual model using the image encoder in CLIP, was proposed by Li et al. [21] for a better visual representation. Additionally, there have been studies [22, 23] focusing on pre-training methods that aim to overcome the domain gap in ReID tasks.
Furthermore, several approaches [24, 25, 26, 27] have utilized extra cues, such as human pose and body parts, to address pose variations and occlusion issues. Similarly, studies [28, 29] have aimed to improve the quality of visual representation through additional identity-guided human semantic parsing and multi-head attention. Nevertheless, alleviating appearance ambiguity is still challenging when relying only on appearance for ReID.
@@ II-B Spatial-temporal ReID
Spatial-temporal information from cameras and target objects has been utilized in many studies to address the constraints of appearance-based ReID. Generally, these studies have taken an appearance-based ReID model as the baseline and leveraged spatial and temporal information. In spatial-temporal ReID, two main challenges occur: 1) estimating spatial-temporal information (the camera network topology) in given camera networks, and 2) utilizing the estimated camera network topology for ReID.
Many studies have attempted to estimate the camera network topology by designing accurate transition time distributions of targets, such as persons and vehicles. For instance, Huang et al. [9] utilized a spatial-temporal model leveraging vehicle pose view embedding and Wang et al. [10] proposed the Histogram–Parzen method to estimate spatial-temporal probability distributions. Additionally, Liu et al. [14, 30] proposed a progressive vehicle ReID that applying simple spatial-temporal information. Similarly, studies [7, 31, 32, 33] have utilized estimated spatial-temporal information to filter out irrelevant gallery images. Furthermore, Shen et al. [8] proposed a Siamese-CNN+Path-LSTM network that predicts the path through visual feature information and spatial-temporal information.
While numerous spatial-temporal ReID methods have been proposed, there are still some limitations. First, the methodologies for estimating spatial-temporal models are simplistic. For example, many methods [7, 14, 30, 31] have built object transition time distributions based on the positive responses between cameras, but noisy and sparse responses make the estimated distributions unreliable. Second, the utilization of spatial-temporal information is not optimized. Some studies [10, 33, 9] have merged both probabilities (i.e., appearance and spatial-temporal) with same importance to obtain the joint probability. Similarly, many methods [7, 31, 32, 33] have applied spatial-temporal information to reduce the search range or perform re-ranking of the initial ReID results.
@@
Figure 3: Examples of estimated transition time distributions between camera pairs. Each bin covers 100 frame ranges. Solid blue lines () mark the estimated distribution ($p_{ij}$) from the histogram ($h_{ij}$) by the proposed adaptive Parzen window (best viewed in color).
III Spatial-temporal Re-identification
based on FusionNet
To address the challenges in ReID, we analyzed the characteristics of objects in camera networks. First, many objects show similar or the same appearances in large-scale camera networks. For example, people can appear similar due to wearing the same clothes (e.g., uniforms and belongings). Especially in the vehicle ReID task, vehicles can look the same according to their model types. Second, the movements of objects between non-overlapping cameras are predictable. E.g., vehicles can only move along roads and highways, rarely deviating from existing roads. Compared to vehicles, people show more complex transition patterns across cameras, but people’s paths are also generally established along the sidewalks and aisles. To summarize, objects show high appearance ambiguities but predictable movements. Therefore, relying only on appearance differences between objects is not effective for the ReID tasks.
Based on these observations, we additionally exploited spatial and temporal relationships between cameras, called a camera network topology. As shown in Fig. 2(a), the proposed spatial-temporal ReID framework consists of two parts: 1) camera network topology estimation based on the adaptive Parzen window, 2) the fusion network on the spatial-temporal probability window. We first built the topology based on the proposed adaptive Parzen window (Section III-A). Then, we trained a FusionNet that optimally combines visual similarity and the camera network topology information for the final ReID prediction (Section III-B).
@@ III-A Adaptive Parzen window
The camera network topology represents the spatial-temporal relationships and connections between cameras, which can be represented by a graph $\mathbf{G}=(\mathbf{V},\mathbf{E})$. The vertices $\mathbf{V}$ represent the cameras, and the edges $\mathbf{E}$ represent the distribution of object transition time. If there are $N_{\text{cam}}$ cameras in the network, the topology is represented as follows:
@@
where $c$ denotes a camera, and $p_{ij}$ denotes the object transition time distribution between camera pairs $c_{i}$ and $c_{j}$.
We build the transition time distribution $p_{ij}$ by leveraging positive pairs from all camera pairs in the training dataset. Using the multiple time differences $(\Delta t)$ of positive pairs, we initially generated a histogram of the transition time $h_{ij}$, which is illustrated by the cyan vertical lines () in Fig. 3. Cho et al. [7] proposed connectivity checking criteria for determining if a pair of cameras are connected by fitting a Gaussian model to the histogram $h_{ij}$. However, this parametric method followed strong assumptions and had difficulty handling outliers and the sparsity of the histogram. Inspired by [10], a Parzen window method can be applied to the initial histograms, and we estimated the probability density function (PDF) of the transition time in a non-parametric manner as follows:
@@
where $\sigma$ is a standard deviation.
Estimating continuous PDFs from discrete histograms efficiently using the Parzen window method is unreasonable when employing a single kernel across diverse histograms from various camera pairs. The strength of the spatial-temporal connection between cameras can be determined by the number of objects passing through those cameras during a certain period [7]. For example, weak connectivity is indicated by few positive pairs between two cameras. However, the Parzen window method extremely enlarges those small responses with a small $\sigma$ value, as depicted in the orange line () in Fig. 3c. In that case, it is better to use a large $\sigma$ value to avoid overfitting the distribution for noise and outliers.
In contrast, if there are many positive pairs between the cameras, the connectivity should be strong. On the other hand, when using a large $\sigma$ value, the resulting distribution becomes uniform, failing to capture any meaningful spatial and temporal relationships between the cameras, as displayed by the green line () in Fig. 3a. In that case, it is better to use a relatively small $\sigma$ value to reflect temporal information between cameras. Therefore, selecting the appropriate $\sigma$ value is important for the quality of the estimated distribution $(p_{ij})$.
To overcome the limitation of the original Parzen window method [10], we newly propose an adaptive Parzen window by setting various $\sigma_{ij}$ values for the camera pairs $(c_{i},c_{j})$. To this end, we designed an adaptive standard deviation according to the different strengths of the camera connectivity as follows:
@@
where $N_{ij}$ denotes the number of positive object pairs between camera $c_{i}$ and $c_{j}$. In addition, $\alpha$ is a scale factor determining the maximum range of $\sigma_{ij}$, and $\beta$ is a smoothness factor that adjusts the sensitivity of $\sigma_{ij}$. The minimum value of $\sigma_{ij}$ cannot be less than 1 unit of the histogram. Then, the values of $\sigma_{ij}$ lie on $\left[1,\alpha\right]$.
By considering the camera indexes, Eq. 2 and Eq. 3 are reformulated as follows:
@@
Therefore, we can estimate reliable distributions ($p_{ij}$) from the initial discrete histograms ($h_{ij}$) by considering the connectivity between cameras. The blue lines () in Fig. 3 are our results based on the adaptive Parzen window.

### v2

@@ Abstract
Object re-identification (ReID) in large camera networks faces numerous challenges. First, the similar appearances of objects degrade ReID performance, a challenge that needs to be addressed by existing appearance-based ReID methods. Second, most ReID studies are performed in laboratory settings and do not consider real-world scenarios. To overcome these challenges, we introduce a novel ReID framework that leverages a spatial-temporal fusion network and causal identity matching (CIM). Our framework estimates camera network topology using a proposed adaptive Parzen window and combines appearance features with spatial-temporal cues within the fusion network. This approach has demonstrated outstanding performance across several datasets, including VeRi776, Vehicle-3I, and Market-1501, achieving up to 99.70% rank-1 accuracy and 95.5% mAP. Furthermore, the proposed CIM approach, which dynamically assigns gallery sets based on camera network topology, has further improved ReID accuracy and robustness in real-world settings, evidenced by a 94.95% mAP and a 95.19% F1 score on the Vehicle-3I dataset. The experimental results support the effectiveness of incorporating spatial-temporal information and CIM for real-world ReID scenarios, regardless of the data domain (e.g., vehicle, person).
@@ I Introduction
An increasing surveillance cameras have been installed in public places for safety, traffic monitoring, and security. However, monitoring all cameras requires significant human effort and resources. Re-identification (ReID) can automatically track targets across multiple non-overlapping cameras, reducing human effort. In general, most studies have focused on the visual appearance of a target to perform ReID. For example, many studies [1, 2] have proposed feature learning methods to represent the appearance of a target. Similarly, to measure feature distance between queries and gallery images effectively, metric learning methods [3, 4] have also been proposed. Recently, ReID performance has been improved by learning visual features and distance metrics with developments in deep learning [5, 6]. These appearance-based methods are robust to pose variations, viewpoint changes, and illumination changes.
(a) Appearance ambiguities
(b) High computational complexities
Figure 1: Challenges of object re-identification in large-scale camera networks due to appearance ambiguities and computational complexities
@@
Nevertheless, appearance ambiguity caused by objects with similar appearances is still not alleviated. As shown in Fig. 1a, vehicles are challenging to match accurately due to their similar appearances, especially when they are of the same model type. In contrast, people generally have more distinct features, but similar appearances can still arise when they wear the same clothing. Additionally, the appearance of many objects across multiple cameras increases computational complexity and reduces identification performance as the number of objects with appearances similar to the target increases (Fig. 1b). Thus, relying only on target appearance is ineffective for the object ReID problem.
@@
Recently, ReID studies using additional spatial and temporal information to alleviate appearance ambiguity [7, 8, 9, 10] have been proposed. Researchers have built a camera-network topology that explains the spatial and temporal relationships between cameras and used the topology to reduce redundant retrieval time ranges for queries. While these methods can potentially improve appearance-based ReID models [5, 6], they still have limitations. Their approaches to camera network topology modeling are too simple, and integrating the appearance model with spatial-temporal information lacks optimization. In addition, most ReID studies have not considered applying ReID in real-world scenarios, such as handling real-time video continuously coming from each camera. However, most ReID studies [5, 11] have performed ReID on fixed and refined query-gallery sets, which are clean and easy to manage. These laboratory settings are difficult to implement in real-world environments and require additional processes such as object appearance management and query-gallery sets management. In summary, previous studies have not fully considered the spatial-temporal dependencies between cameras and the complexities of real-world scenarios.
@@
To consider the spatial and temporal dependencies between cameras, we proposed a spatial-temporal fusion network (FusionNet) in our previous study [12]. For training FusionNet, we proposed a novel adaptive Parzen window method robust to outliers and sparse responses between camera pairs (Section III-A). This method effectively manages varying connection strengths between camera pairs, enabling reliable camera network topology estimation. After estimating the topology, we trained FusionNet to optimally combine appearance similarity and spatial-temporal probabilities (Section III-B). In this study, we extend our study[12] to explore ReID problems in real-world scenarios extensively. To this end, we first built a new large-scale vehicle ReID dataset called Vehicle-3I, which includes 2,038 identities captured by 11 synchronized cameras installed at three different intersections (Section V). Unlike many public datasets [13, 14] that provide only cropped image patches, it provides full-frame videos.
@@
Furthermore, we focus on critical and practical issues for ReID in real-world scenarios. Issue 1): Each identity appears multiple times in the video, necessitating an appearance management process and ID-to-ID ReID. Issue 2): The query and gallery sets cannot be predefined as in previous studies. Instead, they are determined dynamically based on the target objects (query) and their ReID results within the camera network. In Section IV, we propose appearance management and ID-to-ID ReID methods that consider multiple appearances of each identity in the video. Additionally, we implement Causal Identity Matching (CIM), which dynamically adjusts the query and gallery sets according to the camera connection causality.
@@
To evaluate the proposed spatial-temporal fusion network, we tested it on the VeRi776 [14], Vehicle-3I vehicle ReID datasets and the Market-1501 [13] person ReID dataset. We evaluated the effectiveness of the adaptive Parzen window and FusionNet. The proposed spatial-temporal fusion network achieved the best performances across all datasets for rank-1 accuracy (99.70%, 89.22%, 99.11%) and mean average precision (mAP) (91.71%, 83.62%, 93.80%). These results demonstrate that our methods can significantly improve ReID performance across different data domains (e.g., vehicle, person). Additionally, we performed CIM and evaluated the proposed methods in real-world scenarios; the proposed methods achieved superior performances on the Vehicle-3I dataset, with a 94.95% mAP and 95.19% $F_{1}$ score.
@@ II-A Appearance-based ReID
Most ReID studies have focused on developing visual representations of images to distinguish their appearance. Extensive research has been conducted on metric learning and feature learning methods to achieve this objective. A widely studied approach in metric learning involves learning the Mahalanobis distance [3, 4]. Specifically, optimizing triplet loss for deep metric learning [15, 16, 5] has demonstrated outstanding performance in ReID tasks. Ghosh et al. [6] introduced relation-preserving triplet mining (RPTM), a scheme for feature matching-guided triplet mining that ensures triplets preserve natural subgroupings in object IDs.
Ahmed et al. [1] initially employed a deep convolutional neural network (CNN) architecture to capture local relationships between two input images based on mid-level features. Chen et al. [2] improved CNN-based ReID by proposing a deep pyramid feature learning CNN, which can learn scale-specific discriminative features. Similarly, other studies [15, 17] have focused on extracting robust local features. Khamis et al. [18] utilized target attributes for ReID to capture more appearance details. Recently, He et al. [11] proposed TransReID, marking the first attempt to use a transformer to learn robust features from images. Chen et al. [19] utilized the Swin transformer [20] as a backbone for downstream ReID tasks.
Li et al. [21] proposed CLIP-ReID, which fine-tunes the initialized visual model using the image encoder in CLIP for improved visual representation. Additionally, other studies [22, 23] have focused on pre-training methods to overcome the domain gap in ReID tasks. Furthermore, several approaches [24, 25, 26, 27] have utilized extra cues, such as human pose and body parts, to address pose variations and occlusion issues. Similarly, studies [28, 29] have aimed to improve visual representation quality through additional identity-guided human semantic parsing and multi-head attention. Nevertheless, alleviating appearance ambiguity is still challenging when relying only on appearance for ReID.
@@ II-B Spatial-temporal ReID
Many studies have utilized spatial-temporal information from cameras and target objects to address the limitations of appearance-based ReID. Generally, these studies take an appearance-based ReID model as the baseline and leverage spatial and temporal information. In spatial-temporal ReID, two main challenges arise: 1) estimating spatial-temporal information (the camera network topology) in given camera networks, and 2) utilizing the estimated camera network topology for ReID.
Many studies have attempted to estimate the camera network topology by designing accurate transition time distributions of targets, such as persons and vehicles. For instance, Huang et al. [9] utilized a spatial-temporal model leveraging vehicle pose view embedding, and Wang et al. [10] proposed the Histogram-Parzen method to estimate spatial-temporal probability distributions. Additionally, Liu et al. [14, 30] proposed a progressive vehicle ReID that applies simple spatial-temporal information. Similarly, studies [7, 31, 32, 33] have utilized estimated spatial-temporal information to filter out irrelevant gallery images. Furthermore, Shen et al. [8] proposed a Siamese-CNN+Path-LSTM network that predicts paths using both visual and spatial-temporal information.
While numerous spatial-temporal ReID methods have been proposed, limitations remain. First, the methodologies for estimating spatial-temporal models are often simplistic. For example, many methods [7, 14, 30, 31] have built object transition time distributions based on positive responses between cameras, but noisy and sparse responses make the estimated distributions unreliable. Second, the utilization of spatial-temporal information is not optimized. Some studies [10, 33, 9] have merged both probabilities (i.e., appearance and spatial-temporal) with same importance to obtain the joint probability. Similarly, many methods [7, 31, 32, 33] have applied spatial-temporal information to reduce the search range or perform re-ranking of the initial ReID results.
III Spatial-temporal Re-identification
based on FusionNet
To address the challenges in ReID, we analyzed the characteristics of objects in camera networks. First, many objects show similar or the same appearances in large-scale camera networks. For example, people may appear similar because they wear the same clothes (e.g., uniforms and belongings). Especially in the vehicle ReID task, vehicles can look the same according to their model types. Second, the movements of objects between non-overlapping cameras are predictable. For instance, vehicles can only move along roads and highways, rarely deviating from existing routes. Compared to vehicles, people show more complex transition patterns across cameras, but people’s paths are also generally established along the sidewalks and aisles. In summary, objects show high appearance ambiguities but predictable movements. Therefore, relying only on appearance differences between objects is not effective for the ReID tasks.
Based on these observations, we further exploited spatial and temporal relationships between cameras, referred to as camera network topology. As shown in Fig. 2(a), the proposed spatial-temporal ReID framework consists of two parts: 1) camera network topology estimation based on the adaptive Parzen window, 2) the fusion network on the spatial-temporal probability window. We first built the topology based on the proposed adaptive Parzen window (Section III-A). Then, we trained FusionNet to optimally combine visual similarity with camera network topology information for final ReID prediction (Section III-B).
@@ III Spatial-temporal Re-identification
Figure 3: Examples of estimated transition time distributions between camera pairs. Each bin covers 100 frame ranges. Solid blue lines () mark the estimated distribution ($p_{ij}$) from the histogram ($h_{ij}$) using the proposed adaptive Parzen window (best viewed in color).
@@ III-A Adaptive Parzen window
The camera network topology represents the spatial-temporal relationships and connections between cameras, which can be represented as a graph $\mathbf{G}=(\mathbf{V},\mathbf{E})$. The vertices $\mathbf{V}$ represent the cameras, and the edges $\mathbf{E}$ represent the distribution of object transition times. If there are $N_{\text{cam}}$ cameras in the network, the topology is represented as follows:
@@
where $c$ denotes a camera, and $p_{ij}$ denotes the object transition time distribution between the camera pair $c_{i}$ and $c_{j}$.
We build the transition time distribution $p_{ij}$ by leveraging the training dataset’s positive pairs from all camera pairs. Using the multiple time differences $(\Delta t)$ of positive pairs, we initially generated a histogram of the transition time $h_{ij}$, which is illustrated by the cyan vertical lines () in Fig. 3. Cho et al. [7] proposed connectivity checking criteria for determining if a pair of cameras are connected by fitting a Gaussian model to the histogram $h_{ij}$. However, this parametric method relied on strong assumptions and struggled to handle outliers and the sparsity of the histogram. Inspired by [10], we applied a Parzen window method to the initial histograms and estimated the probability density function (PDF) of the transition times in a non-parametric manner as follows:
@@
where $\sigma$ is the standard deviation.
Estimating continuous PDFs from discrete histograms efficiently using the Parzen window method is unreasonable when employing a single kernel across diverse histograms from various camera pairs. The strength of the spatial-temporal connection between cameras can be determined by the number of objects passing through those cameras during a certain period [7]. For example, weak connectivity is indicated by few positive pairs between two cameras. However, the Parzen window method can excessively amplify small responses when using a small $\sigma$ value, as depicted by the orange line () in Fig. 3c. In that case, it is better to use a large $\sigma$ value to avoid overfitting the distribution to noise and outliers.
In contrast, if there are many positive pairs between cameras, the connectivity should be strong. However, when using a large $\sigma$ value, the resulting distribution becomes uniform, failing to capture meaningful spatial and temporal relationships between the cameras, as illustrated by the green line () in Fig. 3a. In such cases, it is better to use a relatively small $\sigma$ value to reflect temporal information between cameras. Therefore, selecting the appropriate $\sigma$ value is important for the quality of the estimated distribution $(p_{ij})$.
To overcome the limitations of the original Parzen window method [10], we newly propose an adaptive Parzen window by setting various $\sigma_{ij}$ values for the camera pairs $(c_{i},c_{j})$. To this end, we designed an adaptive standard deviation according to the different strengths of camera connectivity as follows:
@@
where $N_{ij}$ denotes the number of positive object pairs between cameras $c_{i}$ and $c_{j}$. Additionally, $\alpha$ is a scale factor determining the maximum range of $\sigma_{ij}$, and $\beta$ is a smoothness factor adjusting the sensitivity of $\sigma_{ij}$. The minimum value of $\sigma_{ij}$ cannot be less than 1 unit of the histogram. Therefore, the values of $\sigma_{ij}$ lie within the range of $\left[1,\alpha\right]$.
Considering the camera indices, Eq. 2 and Eq. 3 are reformulated as follows:
@@
Therefore, we can estimate reliable distributions ($p_{ij}$) from the initial discrete histograms ($h_{ij}$) by considering the connectivity between cameras. The blue lines () in Fig. 3 represent our results based on the adaptive Parzen window.